OpenStudy (anonymous):

For what two values of x is the sum of these rational expressions undefined? Type your answer as two integers separated by a comma, like this: -13, 4 x + 3x2– 2x – 15+3x + 12x + 6

OpenStudy (anonymous):

\[\frac{ x+3 }{ x^2-2x-15 }+\frac{ 3x+1 }{ 2x+6 }\]

OpenStudy (anonymous):

@bohotness

OpenStudy (misty1212):

HI!! are you supposed to add these?

OpenStudy (bohotness):

o.o

OpenStudy (anonymous):

yes but you factor them, and I am confused on it.

OpenStudy (misty1212):

the first denominator factors as \((x-5)(x+3)\) the second one as \(2(x+3)\) making the common denominator \[2(x+3)(x-5)\]

OpenStudy (bohotness):

okay i think i might be able to help

OpenStudy (anonymous):

okay, so now i would just add the numerators correct?

OpenStudy (misty1212):

so \[\frac{ x+3 }{ x^2-2x-15 }+\frac{ 3x+1 }{ 2x+6 }\] \[=\frac{(x+3)}{(x-5)(x+3)}\times \frac{2}{2}+\frac{3x+1}{2(x+3)}\times \frac{x-5}{x-5}\]

OpenStudy (anonymous):

im not getting it

OpenStudy (misty1212):

when you add fractions you do not just add up the numerators

OpenStudy (misty1212):

if you added \(\frac{5}{12}+\frac{7}{18}\) what would you do?

OpenStudy (anonymous):

find what would make the same denominator

OpenStudy (misty1212):

yes ok lets agree that that is 36 then what?

OpenStudy (anonymous):

I just dont get the numerator part of what you put up there.

OpenStudy (anonymous):

the numerators would then be 15 and 14

OpenStudy (misty1212):

right you had to multiply by the missing factor when you put both of those over the LCD

OpenStudy (misty1212):

you have to do exactly the same thing with the variables

OpenStudy (anonymous):

so the top is \[3x^2+2x-2\]

OpenStudy (misty1212):

idk i didn't do it you gotta add that up, multiply out and combine like terms